Angle theta is 60o, and F is 40 N. If a 100 kg block is moved from rest along the surface for 100 m, and the surface is frictionless, what is the kinetic energy of the block? What is the final velocity of the block?
Answer #1

here we have

horizontal component of force = f cos 60

= 40 cos 60 = 20 N

acceleration of block a = F/m = 20/100=0.2 m/s^2

initial velocity =u= 0 m/s

let final velocity be v

so , we have v= u + 2as

here s= 100 m

so v= 0+ 2 (0.2)(100) =40 m/s

and kinetic energy = 0.5 (m) (v)2= 0.5 (100) (40)2= 80000 joules